OpenStudy (anonymous):

Find the 20th term and the sum of the first 14 terms of this progression, 2, 5, 8, 11,.........?

OpenStudy (kropot72):

This is an arithmetic progression with the first term a = 2 and the common difference d = 3. The formula for finding the nth term of an AP is \[n _{th}\ term = a+(n-1)d\] To find the 20th term plug the values of a (2), n (20) and d (3) into the formula,

OpenStudy (anonymous):

Given progression, 2, 5, 8, 11,......... is a arithmetic progression Here First term (a) = 2 and Common difference (d) = 5 - 2 = 3 We know, nth term can be represented as, an = a + (n - 1)d and sum of n terms = n/2{2a + (n - 1)d} To find 20th term of the series => a20 = a + 19 d = 2 + 19 * 3 = 2 + 57 = 59 => 20th term = 59 Sum of first 14 terms = 14/2{2 * 2 + (14 - 1) * 3} = 7{4 + 39} = 7 * 43 = 301 => Sum of first 14 terms = 301.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."